TABLE 9.1 Exchange Rates of Major Currencies Currency per U.S. S U.S. $ Currency per U.S. S U.S. Country/Currency EquivalentUse the following table to complete assignment

Suppose that on March 1 of the current year, the peso-US$ exchange rate was P5/$. On March 31 of the current year, the exchange rate stood at P8/$. Calculate the 1-month percent change in the value of the Mexican peso (= P).

Calculate the spot Korean won-Japanese yen exchange rate in W/¥. (Korean won = W; Japanese yen = ¥)

Calculate the spot Taiwanese dollar-euro exchange rate in T$/€. (Taiwanese dollar = T$; euro = €)

Mr. Smith buys 1 share of Canada-based CC Forest Corp. at C$20 (C$ = Canadian currency) when the FX rate is C$1/$. One year later, the price is C$24 and no dividends have been paid. If the current exchange rate is C$1.2/$, calculate the 1-year investment return in U.S. dollars.

Suppose that in the beginning of the year, $1 buys 80 yen (= ¥). The expected annual inflation rates are 6% in Japan and 3% in the U.S., respectively. Calculate the yen-dollar (= ¥/$) exchange rate at the end of the year.

Caterpillar (CAT) of the U.S. just sold an A$1 million crane to Rio Tinto of Australia (currency = A$). Under the terms of the contract, the American company will receive A$1 million in 30 days. Advise Caterpillar on its strategies in dealing with this cash flow. For each strategy, explain how it works and the outcome.

Referring to the same table (Assignment 5 Table), calculate the 1-month Swiss francUK pound forward rate in SF/L and in L/SF. (Swiss franc = SF; pound = L)
Answer #1

1) Mexican peso was P5/$ on March 1 and it is now P8/$ on March 31.

( 8 - 5 ) / 5 = 3 / 5
= 0.6

The change is 60%

2) USD equivalent South Korea Won = 0.0010008
Japanese Yen per USD = 106.81

Won / Yen = 0.0010008 * 106.81
Won / Yen = 0.1069

3) Taiwan dollar / USD = 30.395
USD / Euro = 1.5737

T$ / Euro = 30.395 * 1.5737
T$ / Euro = 47.8326

4) Mr. Smith bought shares of a Canadian company at C$20 and at that time the exchange rate was C$1/$.
Now the stock has rose to C$24 which means it has appreciated by 20%

( 24 - 20 ) / 20 = 0.2 or 20%

On the other hand, the exchange rate also rose to C$1.2/$ which means the currency has depreciated by 20%.
Mr Smith will convert C$ to USD at the new exchange rate and that he will receive $20.

C$24 / 1.2 = $20

The return on investment is zero over the year.

First 4 questions have been answered according to the HOMEWORKLIB RULES.